# AI-based Programming Assistants for Privacy-related Code Generation: The Developers’ Experience

### v1

@@ IV-A2 Who Used AI-based Programming Assistants the Most and Who Used Them the Least
AI code gen tools should provide more user control over the level of compliance with specific privacy regulations, kind of code generated, use of APIs in the code, etc [32]. The developers said that AI-based programming assistants should generate code that is abstract, extensible, and also sample code.
“Generate not that specific code, and then the programmer can extend it” – P44 (Developer, front-end, 6 years exp.)
IV-D8 Recommendations for Data Protection (17.65%; n=9)
@@
AI-based programming assistants need to be able to detect that user-sensitive data are being processed within the target code and be able to handle them properly [26]. Our developers said that the AI code gens should notify and warn the user (i.e., the developer) about possible risks and provide recommendations to the users about ensuring data protection. They further said that AI-based programming assistants should introduce privacy verification checks. But, one developer said that human errors are the main problem source, so the developers need to be well aware of data privacy requirements.
“Well, it could detect that user sensitive data are being processed with the code and ensure that they are handled properly (e.g. inserting into the database). It could also notify the programmer and give him/her recommendations about how to use the generated code to ensure sensitive data safety” – P34 (Developer, mobile, 10 years exp.)
IV-D9 Improve Interpretability (3.92%; n=2)
As an UI improvement, a few developers suggested improving interpretability including visual aspects such as highlighting areas where privacy requirements are violated in generated code.
@@
“The tools (<<<AI code gens>>>) can highlight areas that it thinks can violate privacy requirements. Otherwise, it should be the duty of the programmer to ensure privacy requirements are met” – P41 (Developer, full-stack, 30 years exp.)
V Discussion
@@
The summary of the key challenges and suggestions for improvement of AI-based programming assistants is given in Fig. 11. Our surveyed developers said that AI-based programming assistants are still not mature enough to be used extensively for privacy-preserving code. AI-based programming assistants do not generate code compliant with all privacy requirements and laws. Therefore, relying completely on AI-based programming assistants to generate privacy-related code is not a wise decision. Our developers had concerns about using AI-based programming assistants because of the distrust they have regarding both their training data and their provided code context data. Developers want more information about generated code explanation, source, and compliance to different privacy-related requirements. They recommend not to use real data when generating code for the app in production, but rather use examples to generate code. Key priorities for improvement include improving model training and introducing robust algorithms that understand data privacy laws, self-testing the generated code, generating explainable outputs, improving transparency, following standards, having compliance, improving ease of use, recommendations for privacy-preserving requirements, and improving interpretability.
Figure 11: The summary of the key challenges and suggestions for improvement of AI-based programming assistants.
V-A Implications for Practice
V-A1 Model and system-level improvements
@@
Better model training and introducing robust algorithms that understand multiple data privacy laws, code that follows standards, generating more explainable output, and running self-tests before providing code to developers. Calibrating the AI-based programming assistants against industry-standard trustworthy AI checks will help mitigate unnecessary threats and disadvantages for AI-based programming assistants to survive in the market.
V-A2 Human-AI interaction improvements
@@
Ease of use in prompting and including generated code, interpretability, highlighting sensitive data, highlighting non-compliant code, and notifying/providing recommendations for privacy approaches can enhance human-AI interaction.
V-A3 Organisational improvements
@@
Organisations owning AI-based programming assistants need to provide better compliance and improve transparency about their models and how data is gathered, processed and models trained. This is extremely important to gain trust from users, both code developers and code end users.
V-B Limitations and Implications for Research
AI-based programming assistants might have changed in various ways, hence the experience with them could be different from what our participants experienced, and what we experienced. Therefore, as a next step, researchers may continue conducting research on these tools/ replicating our study to see the progress of the AI-based programming assistants (if they have made any).
How the AI-based programming assistants behave may differ from case to case. In this paper, we have focused on generating code for privacy. However, the results may be different in other cases, for example when generating code for other functional/non-functional requirements. In the future, researchers may consider conducting research to see how AI-based programming assistants behave across various use cases.
The number of participants who participated in our study was limited, and we did not have the chance to ask follow-up questions. To gain a more in-depth understanding of the developers’ experience with AI-based programming assistants, further research can be conducted using other data collection methods such as interviews.
Conducting experiments with professional developers will help in further understanding their expectations and their interactions with AI-based programming assistants. In future, experiments with professional developers can be conducted to better understand human (developer)-AI interaction.
V-C Related Work
Even though the experiment the authors of [9] ran were not related to privacy, they used medium and hard problems on Leetcode. The authors found that ChatGPT struggles to generate code for new and unseen problems. [33] states that Stack Overflow slightly outperforms ChatGPT when answering privacy questions, but ChatGPT generated code can be used as alternatives. [34] found cases where insecure code was generated by GitHub Copilot. According to [35], when GitHub Copilot generated large chunks of code, the participants in their study stated that the code was cumbersome and it required significant effort to understand the meaning of generated code. As per [14], the first-time users of GitHub Copilot spend more time reviewing the code generated by the AI code generator than writing code.The authors also mention in their work that he early users of GitHub Copilot reported that there was a risk of revealing secrets such as API keys and passwords or suggest inappropriate test. Further, the authors state that this was fixed at the technical preview of GitHub Copilot. However, according to our findings, the distrust of AI-based programming assistants and use of alternatives to ensure privacy still exist. [36] mentions that developers’ programming shifts from writing code to checking and doing unfamiliar debugging.
@@ V-C Related Work
[37] recommends that it is necessary to assure users about privacy and confidentiality. According to what we found, improving transparency will help in that case. [34] touch upon the fact that GitHub Copilot is a black-box and a closed resource residing on a remote server which is unaccessible by general users. [37] raises that getting user content is key as per GDPR privacy principles. The authors recommend the users of AI code generating tools need to get the informed consent from paricipants. This is also related to being transparent (the previous point) as in order to get user consent, the tool needs to be transparent about its actions. As per [9], long prompts tend to have a negative impact on code generated by ChatGPT and Codex. Therefore, the participants need to master prompt engineering. [13] found that longer details result in poor outputs. By going through experience reports of developers, the authors of [36] found that the developers find it hard to write prompts. [37] recommends that AI code generating tools have to warn users in cases like sharing information to the public. [35] states the participants found that the code generated in general is similar to style and quality of a human but for a couple of the experiments the participants participated found that the code generated were not similar to the human style. As mentioned previously, they also found the generation of large chunks of code cumbersome.
To the best of our knowledge, we could not find any particular study specifically focusing on AI-based code generation for privacy that revealed insights on the need of understanding diverse data privacy laws, self testing, explainable output, compliance which are imperative findings of our study.
VI Conclusion
@@
AI-based programming assistants have rapidly become very popular. We conducted a survey of 51 professional developers that identified a range of using, perceived usefulness, and concerns about these tools. We then analysed the literature and conducted a targeted study using ChatGPT to generate code that should have GDPR-compliant privacy-preserving features. Our study revealed common usage challenges and improvements for AI-based programming assistants available in the market. AI-based programming assistants do need to be improved more for developers to trust them with generating code for ensuring data protection (privacy). We developed several recommendations for developers and AI-based programming assistants producing companies to improve them for privacy-preserving code generation tasks.
Acknowledgment

### v2

@@ 4.3. RQ2 – Key Challenges in Generating Code Ensuring Privacy using AI-based Programming Assistants
AI code gen tools should provide more user control over the level of compliance with specific privacy regulations, kind of code generated, use of APIs in the code, etc (Hussain et al., 2020). The developers said that AI-based programming assistants should generate code that is abstract, extensible, and also sample code: “Generate not that specific code, and then the programmer can extend it” – P44 (Developer, front-end, 6 years exp.)
4.4.8. Recommendations for Data Protection (17.65%; n=9)
@@
AI-based programming assistants need to be able to detect that user-sensitive data are being processed within the target code and be able to handle them properly (Gupta et al., 2023). Our developers said that the AI code gens should notify and warn the user (i.e., the developer) about possible risks and provide recommendations to the users about ensuring data protection. They further said that AI-based programming assistants should introduce privacy verification checks. But, one developer said that human errors are the main problem source, so the developers need to be well aware of data privacy requirements: “Well, it could detect that user sensitive data are being processed with the code and ensure that they are handled properly (e.g. inserting into the database). It could also notify the programmer and give him/her recommendations about how to use the generated code to ensure sensitive data safety” – P34 (Developer, mobile, 10 years exp.)
4.4.9. Improve Interpretability (3.92%; n=2)
@@
As an UI improvement, a few developers suggested improving interpretability including visual aspects such as highlighting areas where privacy requirements are violated in generated code: “The tools (<<<AI code gens>>>) can highlight areas that it thinks can violate privacy requirements. Otherwise, it should be the duty of the programmer to ensure privacy requirements are met” – P41 (Developer, full-stack, 30 years exp.)
5. Discussion
@@
Our surveyed developers said that AI-based programming assistants are still not mature enough to be used extensively for privacy-preserving code. AI-based programming assistants do not generate code compliant with all privacy requirements and laws. Therefore, relying completely on AI-based programming assistants to generate privacy-related code is not a wise decision. Our developers had concerns about using AI-based programming assistants because of the distrust they have regarding both their training data and their provided code context data. Developers want more information about generated code explanation, source, and compliance to different privacy-related requirements. They recommend not to use real data when generating code for the app in production, but rather use examples to generate code. Key priorities for improvement include improving model training and introducing robust algorithms that understand data privacy laws, self-testing the generated code, generating explainable outputs, improving transparency, following standards, having compliance, improving ease of use, recommendations for privacy-preserving requirements, and improving interpretability.
5.1. Implications for Practice
@@
Model and system-level improvements. Better model training and introducing robust algorithms that understand multiple data privacy laws, code that follows standards, generating more explainable output, and running self-tests before providing code to developers. Calibrating the AI-based programming assistants against industry-standard trustworthy AI checks will help mitigate unnecessary threats and disadvantages for AI-based programming assistants to survive in the market.
@@
Human-AI interaction improvements. Ease of use in prompting and including generated code, interpretability, highlighting sensitive data, highlighting non-compliant code, and notifying/providing recommendations for privacy approaches can enhance human-AI interaction.
@@
Organisational improvements. Organisations owning AI-based programming assistants need to provide better compliance and improve transparency about their models and how data is gathered, processed and models trained. This is extremely important to gain trust from users, both code developers and code end users.
5.2. Limitations and Implications for Research
AI-based programming assistants might have changed in various ways, hence the experience with them could be different from what our participants experienced, and what we experienced. Therefore, as a next step, researchers may continue conducting research on these tools/ replicating our study to see the progress of the AI-based programming assistants (if they have made any). How the AI-based programming assistants behave may differ from case to case. In this paper, we have focused on generating code for privacy. However, the results may be different in other cases, for example when generating code for other functional/non-functional requirements. In the future, researchers may consider conducting research to see how AI-based programming assistants behave across various use cases. The number of participants who participated in our study was limited, and we did not have the chance to ask follow-up questions. To gain a more in-depth understanding of the developers’ experience with AI-based programming assistants, further research can be conducted using other data collection methods such as interviews. Conducting experiments with professional developers will help in further understanding their expectations and their interactions with AI-based programming assistants. In future, experiments with professional developers can be conducted to better understand human (developer)-AI interaction.
5.3. Related Work
@@
Even though the experiment conducted by Tian et al. (Tian et al., 2023) was not related to privacy, they used medium and hard problems on Leetcode. The authors found that ChatGPT struggles to generate code for new and unseen problems. Delile et al. (Delile et al., 2023) state that Stack Overflow slightly outperforms ChatGPT when answering privacy questions, but ChatGPT generated code can be used as alternatives. Pearce et al. (Pearce et al., 2022) found cases where insecure code was generated by GitHub Copilot. According to Tang et al. (Tang et al., 2023), when GitHub Copilot generated large chunks of code, the participants in their study stated that the code was cumbersome and it required significant effort to understand the meaning of generated code. As per Bird et al.’s work (Bird et al., 2023), the first-time users of GitHub Copilot spend more time reviewing the code generated by the AI code generator than writing code.The authors also mention in their work that he early users of GitHub Copilot reported that there was a risk of revealing secrets such as API keys and passwords or suggest inappropriate test. Further, the authors state that this was fixed at the technical preview of GitHub Copilot. However, according to our findings, the distrust of AI-based programming assistants and use of alternatives to ensure privacy still exist. Sarkar et al.’s work (Sarkar et al., 2022) mentions that developers’ programming shifts from writing code to checking and doing unfamiliar debugging. Cheng et al. (Cheng et al., 2022) recommend that it is necessary to assure users about privacy and confidentiality. According to what we found, improving transparency will help in that case. Pearce et al. (Pearce et al., 2022) touch upon the fact that GitHub Copilot is a black-box and a closed resource residing on a remote server which is unaccessible by general users. Cheng et al. (Cheng et al., 2022) raises that getting user content is key as per GDPR privacy principles. The authors recommend the users of AI code generating tools need to get the informed consent from paricipants. This is also related to being transparent (the previous point) as in order to get user consent, the tool needs to be transparent about its actions. As per the research by Tian et al. (Tian et al., 2023), long prompts tend to have a negative impact on code generated by ChatGPT and Codex. Therefore, the participants need to master prompt engineering. Camara et al. (Cámara et al., 2023) found that longer details result in poor outputs. By going through experience reports of developers, the authors of (Sarkar et al., 2022) found that the developers find it hard to write prompts. Cheng et al. (Cheng et al., 2022) recommend that AI code generating tools have to warn users in cases like sharing information to the public. In their work (Tang et al., 2023), Tang et al. state the participants found that the code generated in general is similar to style and quality of a human but for a couple of the experiments the participants participated found that the code generated were not similar to the human style. As mentioned previously, they also found the generation of large chunks of code cumbersome. To the best of our knowledge, we could not find any particular study specifically focusing on AI-based code generation for privacy that revealed insights on the need of understanding diverse data privacy laws, self testing, explainable output, compliance which are imperative findings of our study.
6. Conclusion
@@
AI-based programming assistants have rapidly become very popular. We conducted a survey of 51 professional developers that identified a range of using, perceived usefulness, and concerns about these tools. Our study revealed common usage challenges and improvements for AI-based programming assistants available in the market. AI-based programming assistants do need to be improved more for developers to trust them with generating code for ensuring data protection (privacy). We developed several recommendations for developers and AI-based programming assistants producing companies to improve them for privacy-preserving code generation tasks.
Acknowledgments